Anna has 12 bills in her wallet, some 5$ and some $10. The total value of the bills is $100. How many of each bill does Anna have?
Answer by stanbon(48558)

Anna has 12 bills in her wallet, some 5$ and some $10. The total value of the bills is $100. How many of each bill does Anna have?
-------------
Let # of $5 bills be x ; Value of these is 5x dollars
# of $10 bills is "12-x" ; Value of these is 10(12-x)= 120-10x dollars
---------
EQUATION:
value + value = 100 dollars
5x+120-10x = 100
-5x = -20
x=4 ( # of $5 bills)
12-x=8 (# of $10 bills)
============
Cheers,
Stan H.